OpenStudy (anonymous):

HELP: if 9 is added to twice a number and this sum is multiplied by 4, the result is the same as if the number is multiplied by 4 and 8 is added to the product. What is the number?

OpenStudy (anonymous):

Ans. -7

OpenStudy (anonymous):

\[(9+2x)4=4x+8 \]

OpenStudy (nurali):

(9+2x)*4 = 4x + 8 solving for x 8x + 36 = 4x + 8 4x = -32 x = -8

OpenStudy (anonymous):

"x" is the number

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

My pleasure :)